MOH Enterprise is a firm that specializes in the production of canned tomatoes. The first two 

columns in the following table give the firm’s short-run production function when the only variable 

input is labor, and capital (the fixed input) is held constant at 5 units. The price of capital is Ghc 

2,000 per unit, and the price of labor is Ghc 500 per unit.

L Q APL MPL

Cost Average cost

TFC TVC TC AFC AVC ATC MC

0 0

20 4000

40 10000

60 15000

80 19400

100 23000

i. Complete the table by calculating the two measures of productivity (APL and MPL) and the 

various categories of cost corresponding to each number of workers.

ii. Using the answer in (i) above, what is the relation between average variable cost and 

marginal cost of the firm? 

iii. Using the answer in (i) above, what is the relation between average product and average 

variable cost of the firm?


1
Expert's answer
2021-06-03T18:40:19-0400
  1. )

"APL=\frac {Q}{L}"


"MPL= \frac {\Delta Q}{\Delta L}"


"TFC= total no. of fixed inputs \times price per unit"

"TVC= total no. of variable inputs \times price per unit"

"TC=TFC +TVC"

"AFC=\frac {TFC}{Q}"


"AVC= \frac {TVC}{Q}"


"ATC= \frac {TC}{Q}"


"MC= \frac {\Delta TC}{\Delta Q}"





(2.)AVC and MC increase at different rates at L=0. At L=40 onwards, AVC and MC increase at almost the same rate as L is increased. AVC= MC at L=60

(3.)Average product increases up to 250 where L =40. At L=40 and L=60, average product begins to fall as L increases. AVC increases up to where L=40, thereafter as L increases, AVC does not increase, its value remains constant.